Find the eccentricity of the ellipse 3x
2 + 4y
2 = 12 and the equation of the tangent to
the ellipse at the point (1, 3/2). If this tangent meets the y-axis at the point G, and
S and S
′ are the foci of the ellipse, find the area of triangle SS′G.
1
Expert's answer
2020-03-05T12:20:18-0500

"3x^2+4y^2=12"

divided by 12 on both sides

"{3 \above{2pt} 12}x^2+{4 \above{2pt} 12}y^2=1"

"{x^2 \above{2pt} 4}+{y^2 \above{2pt} 3}=1"

The eccentricity is a measure of how much the ellipse deviates from the circle.For an Ellipse with major axis parallel to the x-axis, the eccentricity is "{ \sqrt{a^2-b^2}\above{2pt} a}"

"=" "{ \sqrt{a^2-b^2}\above{2pt} a}"

Calculate Ellipse properties

"3x^2+4y^2=12" : Ellipse with center "(h,k)=(0,0)" , semi-major axis "a=2" ,semi -minor axis

"b=\sqrt{3}"

"3x^2+4y^2=12"

Ellipse Standard Equation:

"{(x-h)^2 \above{2pt} a^2}+{(y-k)^2 \above{2pt} b^2}=1" is the Ellipse standard equation with center (h,k) and a,b are the semi-major and semi-minor axes .

Rewrite "3x^2+4y^2=12" in the form of standard ellipse equation .

"3x^2+4y^2=12"

divided by 12 on both sides

"{3 \above{2pt} 12}x^2+{4 \above{2pt} 12}y^2=1"

"{x^2 \above{2pt} 4}+{y^2 \above{2pt} 3}=1"

Rewrite in standard form

"{(x-0)^2 \above{2pt} 2^2}+{(y-0)^2 \above{2pt} (\sqrt{3})^2}=1"

Therefore Ellipse properties are :

"(h,k)=(0,0), a=2 ,b=\sqrt{3}"

"a>b" therefore a is semi-major axis and b is semi-minor axis.

Ellipse with center "(h,k)=(0,0)" ,semi-major axis "a=2" , semi-minor axis "b=\sqrt{3},"

"{ \sqrt{2^2-(\sqrt{3})^2}\above{2pt} 2}"

"={1 \above{2pt} 2}"



the equation of the tangent to

the ellipse at the point (1, 3/2). If this tangent meets the y-axis at the point G, and

S and S

Equation tangent will be

"y=mx+\sqrt{a^2m^2+b^2}"

"y=mx+\sqrt{4m^2+(\sqrt{3})^2}"

As this line passes through (1,3/2 )

Therefore "{3 \above{2pt} 2}-m=" "\sqrt{4m^2+(\sqrt{3})^2}"

"\implies" "{9 \above{2pt} 4}+m^2-3m=4m^2+(\sqrt{3})^2"

"\implies 9+4m^2-12m=16m^2+4(\sqrt{3})^2"

"\implies 9+4m^2-12m-16m^2-12=0"

"\implies -12m^2-12m-3=0"

by quadratic formula

"m=-{1 \above{2pt} 2}"

"y=-{1\above{2pt} 2}x+2"


Area of triangle SS'G="{1 \above{2pt} 2}\sqrt{({1 \above{2pt} 2})^2+(1)^2}"

"={1 \above{2pt} 2}" "\sqrt{{5 \above{2pt} 4}}"

"={1 \above{2pt} 4}\sqrt{5}"